OpenStudy (rissyroo118):

math questions help!! the ratio 16 : 200 as a fraction in simplified form is?

OpenStudy (rissyroo118):

2/25?

OpenStudy (anonymous):

16/200 = 8/100 = 4/50 = 2/25

OpenStudy (rissyroo118):

cool thanks so much!

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

haha np